Suppose that a firms out put is given by: q = αLK2 Where α is a productivity factor.

i. Specify the marginal products of labour and capital.

ii. If the cost function is C = wL + rK. Specify the quanitites of capital and Labour that minimize costs?


1
Expert's answer
2021-07-18T18:38:01-0400

i.

The marginal product of capital "\frac{∂Q
}{∂L }= αK^2"


The marginal product of labor "\frac{∂Q
}{∂K }= 2αLK"


ii

"C = wL + rK\\"

When plotted with capital on the vertical axis and has intercepted at K =r and L = w.

The cost-minimizing point is a corner solution.